# DPA: Dual Prototypes Alignment for Unsupervised Adaptation of Vision-Language Models

### v1

@@ 4 Experiments
Analysis of Model Components: We investigate the impact of different components through the training of four distinct models: Base: This model iteratively generates pseudo-labels using textual prototypes and incorporates regularization loss through self-training. Center: CLIP undergoes training with PLs generation using image and text prototypes, as detailed in Section 3.2, in addition to building upon the foundation established in the Base model. Center+$w$: introduces the weighting strategy for Center model. Center+$w$+Align: introduces the alignment loss, as detailed in Section 3.2. As shown in Table 2, Base model, which relies solely on self-training, demonstrates a notable improvement compared to the zero-shot CLIP. However, as training progresses, the impact of confirmation bias becomes evident, leading to a decrease in the accuracy of the pseudo-labels, as illustrated in Figure 3-(a). To mitigate the effects of confirmation bias, we introduce the pseudo-label generation strategy in  Center model. This approach enhances the quality of pseudo-labels throughout the training process, leading to a significant gain of $+7.28\%$ over zero-shot CLIP in average accuracy. Incorporating the weighting strategy in Center+$w$ model results in an additional increase of $+8.01\%$ over zero-shot CLIP in accuracy. By assigning appropriate weights to the self-training loss, the model can better balance the contributions of high-confidence and low-confidence pseudo-labels, leading to more stable and practical training. Finally, including the alignment loss in Center+$w$+Align model leads to a further increase of $+8.58\%$ over zero-shot CLIP in average accuracy. By explicitly encouraging the alignment between image and text prototypes, the model learns more robust and transferable representations, which are crucial for effective unsupervised adaptation. Finally, we conduct a detailed analysis of the evolution of pseudo-label accuracy during training epochs for our method (DPA) compared to Base model (Figure 3-(a,b,c)).
Table 2: Ablation study. See text for details.
@@
Figure 3: Base vs. DPA in (a) PLs, (b) training and (c) testing accuracy on EuroSAT dataset.
Incorporating Image Prototypes as Pseudo-Label Source: We analyse the impact of incorporating image prototypes as an additional classifier. We exclude Equation 7 and generate PLs solely based on text prototypes, while preserving the alignment of image-text prototypes and the weighting strategy. We update $\mathbf{MB}$ solely using the pseudo-labels derived from the textual prototypes. Table 3 shows that the alignment and weighting strategies generally perform well across datasets, with minor differences compared to the entire model, except for EuroSAT and Flowers-102.
@@
Impact of Image Prototype Initialization on Performance: We evaluate three initializations for image prototype methods: mean, weighted mean, and similarity weighted mean [28] (Table 4). The results demonstrate significant improvements over the baselines for all initialization strategies.
@@
Robustness Analysis under Distribution Shifts: To evaluate robustness against natural distribution shifts, we show results on ImageNet-Sketch [57], ImageNet-Adversarial [58], and ImageNet-Rendition [59]. Initially, we evaluate a fine-tuned ViT-B/32 model on ImageNet using our method, achieving an accuracy of $64.64\%$ on the ImageNet validation set. We explore transductive transfer learning [60, 61] with our approach. Transductive learning involves assuming access to unlabelled test images from the new distribution (Table 5). We observe a modest enhancement in performance under distribution shift in 2 out of 3 datasets. This slight improvement is attributed to utilizing only $25\%$ of ImageNet dataset, a deliberate choice made to manage computational resources effectively and address the substantial size of ImageNet training dataset.
Table 5: Transductive adaptation to distribution shift
Method ImageNet ImageNet-S ImageNet-A ImageNet-R
@@
Full Fine-tuning: We compare full fine-tuning of the models instead of solely fine-tuning the layer-normalization weights (Table 6) under noisy PLs. The last row shows the PLs of zero-shot CLIP used at the start of the training for constructing the initial image prototypes. In datasets with noisy PLs, Table 6 highlights the importance of solely fine-tuning layer-normalization weights during adaptation to prevent overfitting and maximize performance.
Table 6: Comparison of full fine-tuning vs. layer-normalization weights fine-tuning for DPA
Zero-shot CLIP 90.69 44.42 43.84 19.50 82.40 66.46 87.50 58.74 89.80 65.10 64.20 64.46
DPA (LayerNorm Tuning) 96.06 55.69 80.04 20.67 84.76 75.56 90.71 62.62 95.97 76.47 68.49 73.31
DPA (Full Tuning) 94.57 53.03 62.92 17.37 82.62 75.23 91.14 54.06 97.44 78.91 68.57 70.53
DPA (PLs accuracy of Zero-shot CLIP) 88.83 43.54 44.90 17.79 77.91 66.53 83.80 57.81 89.64 65.60 64.10 63.68
With other VLMs: We evaluate the efficacy of DPA with other VLM models in Table 7. To assess this, we conducted experiments on SLIP [62], a version of CLIP incorporating self-supervised learning objectives during pre-training. Additionally, we evaluated DPA using a different CLIP-based architecture, rather than ViT-B/32. As depicted in Table 7, DPA consistently demonstrates substantial and consistent improvements across diverse VLMs and architectures.
Table 7: Effectiveness of DPA on different model architectures and pre-training strategies.
Method Caltech DTD ESAT FGVCA Food Flower OxPets StCars CIFAR10 CIFAR100 UCF
Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt Init $\rightarrow$ Adapt
SLIP (ViT-L/16) 83.90 $\rightarrow$ 90.20 25.53$\rightarrow$ 36.28 20.90$\rightarrow$ 31.50 8.40$\rightarrow$ 9.18 60.45$\rightarrow$ 68.08 64.23$\rightarrow$ 75.07 33.22$\rightarrow$ 44.02 8.00$\rightarrow$ 9.03 81.67$\rightarrow$ 88.10 48.70$\rightarrow$ 59.88 38.09$\rightarrow$ 52.76
DPA (ViT-B/16) 92.60$\rightarrow$ 96.09 44.70$\rightarrow$ 50.69 49.00$\rightarrow$ 81.22 23.97$\rightarrow$ 25.14 87.80$\rightarrow$ 89.83 70.89$\rightarrow$ 78.68 89.00$\rightarrow$ 93.19 64.70$\rightarrow$ 68.45 90.80$\rightarrow$ 96.43 68.22$\rightarrow$ 78.10 69.10 $\rightarrow$ 74.62
Comparison with ReCLIP: We provide a comparison with a very recent method for unsupervised adaptation of VLMs, namely ReCLIP [16] (Table 8). DPA consistently outperforms ReCLIP [16] (Table 8) in both inductive and transductive settings. Also, DPA has superior training efficiency and performance with fewer parameters and reduced computational complexity (Figure 4).
@@
Comparison with Few-Shot methods: We compare with the few-shot methods (CoOp [7] and MaPLe [9]) in Table 9. DPA match or surpass few-shot methods in 6 of 11 datasets.
Table 9: Comparison with few-shot methods across multiple datasets.
@@
Limitations and broader impact: While our method shows promising results, there is still room for improvement in mitigating zero-shot CLIP’s strong confirmation biases, as illustrated in Figure 5, which was not the primary focus of this work. This figure underscores the remarkable efficacy of our approach in source-free adaptation using dual classifiers to generate pseudo-labels to reduce false positive rates across all categories. However, zero-shot CLIP has strong confirmation biases, resulting in our method consistently achieving false positive predictions for certain classes. The confusion between classes is also evident in ReCLIP, indicating that addressing these strong biases requires more robust de-biasing strategies.
((a)) Zero-shot CLIP
((b)) ReCLIP (inductive training)
((c)) DPA (Ours)
Figure 5: Confusion matrices of zero-shot CLIP, ReCLIP [16], and our method (DPA).
@@ 5 Conclusion
We present DPA, an unsupervised domain adaptation method for VLMs that aims at bridging the domain gap between visual and textual representations. DPA introduces a novel idea of dual prototypes, which work as two distinct classifiers, and their outputs are fused via a convex combination. It also ranks pseudo-labels for robust self-training during early training. DPA also enhances the alignment of visual-textual prototypes to improve adaptation performance further. DPA shows improved pseudo-labelling accuracy, which leads to notable performance improvements on various downstream tasks. DPA significantly outperforms zero-shot CLIP and the state-of-the-art baselines across 13 downstream vision tasks.

### v2

@@ 4 Experiments
Sensitivity Analyses of Hyperparameters: The sensitivity analyses of the hyperparameters $\lambda_{2}$, $\lambda_{3}$, and $\beta$ on the ESAT dataset are presented in Fig. 3.
[…]
Results on Different Backbones We compare different backbones (ViT-B/16) in Tab. 2  [^1] A fair comparison using RN50 is not possible, as our relevant baselines were specifically designed for the ViT image encoder architecture (For example, LaFTer [33] / ReCLIP [17]).
Method Caltech DTD ESAT FGVCA Flower OxPets StCars CIFAR10 CIFAR100 UCF Avg
Zero-shot CLIP 92.60 44.70 49.00 23.97 70.89 89.00 64.70 90.80 68.22 69.10 66.30
UPL 95.14 45.90 55.36 22.53 73.93 87.98 60.33 92.92 70.41 67.43 67.19
POUF 95.40 48.60 59.50 24.40 72.10 91.80 63.50 92.30 66.30 71.50 68.54
LaFTer 95.92 54.79 72.10 22.38 75.15 85.28 64.72 96.50 76.30 67.20 71.03
DPA 96.09 50.69 81.22 25.14 78.68 93.19 68.45 96.43 78.10 74.62 74.26
Table 2: Top-1 accuracy (%) comparison of DPA with SOTA unsupervised adaptation methods using CLIP-ViT-B/16.
Analyses of Model Components: We investigate the impact of different components through the training of four distinct models: Base: This model iteratively generates pseudo-labels using textual prototypes and incorporates regularization loss through self-training. Center: CLIP undergoes training with PLs generation using image and text prototypes, as detailed in Section 3.2, in addition to building upon the foundation established in the Base model. Center+$w$: introduces the weighting strategy for Center model. Center+$w$+Align: introduces the alignment loss, as detailed in Section 3.2. As shown in Tab. 3, Base model, which relies solely on self-training, demonstrates a notable improvement compared to the zero-shot CLIP. However, as training progresses, the impact of confirmation bias becomes evident, leading to a decrease in the accuracy of the pseudo-labels, as illustrated in Fig. 4-(a). To mitigate the effects of confirmation bias, we introduce the pseudo-label generation strategy in the  Center model. This approach enhances the quality of pseudo-labels throughout the training process, leading to a significant gain of $+7.28\%$ over zero-shot CLIP in average accuracy. Incorporating the weighting strategy in Center+$w$ model results in an additional increase of $+8.01\%$ over zero-shot CLIP in accuracy. By assigning appropriate weights to the self-training loss, the model can better balance the contributions of high-confidence and low-confidence pseudo-labels, leading to more stable and practical training. Finally, including the alignment loss in Center+$w$+Align model leads to a further increase of $+8.58\%$ over zero-shot CLIP in average accuracy. By explicitly encouraging the alignment between image and text prototypes, the model learns more robust and transferable representations, which are crucial for effective unsupervised adaptation. Finally, we show a detailed analysis of the evolution of pseudo-label accuracy during training epochs for our method (DPA) compared to Base model in Fig. 4-a. See more discussions in suppl.
@@
Table 3: Ablation study. See text for details.
Figure 4: Base vs. DPA in (a) PLs, (b) training, and (c) testing accuracy on EuroSAT [12] dataset.
Incorporating Image Prototypes as Pseudo-Label Source: We analyze incorporating image prototypes as an additional classifier. We exclude  Eq. 7 and generate PLs solely based on text prototypes while preserving the alignment of image-text prototypes and the weighting strategy. We update $\mathbf{MB}$ solely using the pseudo-labels derived from the textual prototypes.  Table 4 shows that the alignment and weighting strategies generally perform well across datasets, with minor differences compared to the entire model, except for EuroSAT and Flowers-102.
@@
Impact of Image Prototype Initialization: We evaluate three initializations for image prototype methods: mean, weighted mean, and similarity weighted mean [28] (Table 5). The results demonstrate significant improvements over the baselines for all initialization strategies.
@@
Robustness Analysis under Distribution Shifts: To evaluate robustness against natural distribution shifts, we show results on ImageNet-Sketch [47], ImageNet-Adversarial [14], and ImageNet-Rendition [13]. Initially, we evaluate a fine-tuned ViT-B/32 model on ImageNet using our method, achieving an accuracy of $64.64\%$ on the ImageNet validation set. We explore transductive transfer learning [51, 39] with our approach. Transductive learning involves assuming access to unlabeled test images from the new distribution (Table 6).
Method ImageNet ImageNet-S ImageNet-R ImageNet-A
@@
Table 6: Transductive adaptation to distribution shift.
Evaluation of Fine-Tuning Techniques: Table 7 shows the results of various fine-tuning techniques for CLIP, including LoRA [15], K-Adaptation [11], and full model fine-tuning, and our approach, which involves fine-tuning only the layer normalization weights (see Tab. 7). The last row of the table shows the PLs of zero-shot CLIP used at the start of the training for constructing the initial image prototypes. In datasets characterized by highly noisy PLs, Tab. 7 underscores the significance of exclusively fine-tuning layer-normalization weights during adaptation. This approach helps to mitigate overfitting and enhances overall performance. Our findings indicate that fine-tuning solely the layer normalization weights is effective and stable for adapting models under noisy supervision.
Method Caltech DTD ESAT FGVCA Flower OxPets StCars Avg
Zero-shot CLIP 90.69 44.42 43.84 19.50 66.46 87.50 58.74 58.74
DPA (LoRA [15]) 90.87 50.32 65.16 19.53 66.87 87.71 57.59 62.58
DPA (K-Adaptation [11]) 92.49 52.50 69.80 19.38 68.09 89.04 60.81 64.59
DPA (Full Tuning) 94.57 53.03 62.92 17.37 75.23 91.14 54.06 64.05
DPA (LN) 96.06 55.69 80.04 20.67 75.56 90.71 62.62 68.76
DPA (PLs accuracy of Zero-shot CLIP) 88.83 43.54 44.90 17.79 66.53 83.80 57.81 57.6
Table 7: Top-1 accuracy (%) comparison of DPA using different parameter-efficient fine-tuning methods. All experiments are conducted with CLIP-ViT-B/32 as the backbone.
Data Scale and Downstream Task Performance: We conducted an experiment using the CIFAR-10 dataset to evaluate the scalability of our method (DPA), compared to LaFTer, as illustrated in Fig. 5. To assess the robustness of our approach, we tested various portions of the dataset: 20%, 40%, 60%, and 80%. Our findings indicate that the DPA method consistently performs well across different dataset sizes, demonstrating its scalability regardless of the amount of data used.
[…]
Comparison with ReCLIP: We provide a comparison with a very recent method for unsupervised adaptation of VLMs, namely ReCLIP [17] (Table 8). DPA consistently outperforms ReCLIP [17] in both inductive and transductive settings. Also, DPA has superior training efficiency and performance with fewer parameters and reduced computational complexity comparing the baselines (Figure 6).
@@
Meth. Caltech DTD ESAT FGVCA Food Flower OxPets StCars CIF.10 CIF.100 UCF Avg
ZS 90.69 44.42 43.84 19.50 82.40 66.46 87.50 58.74 89.80 65.10 64.20 64.79
@@
Table 9: Comparison with few-shot methods across 11 different datasets. DPA deliver competitive performance or outperform few-shot methods in 6 of 11 datasets.
Comparison with Few-Shot methods: Table 9 provides a performance comparison of DPA with the few-shot methods (CoOp [61] and MaPLe [21]). DPA show competitive performance or surpasses few-shot methods in 6 of 11 datasets.
@@ 5 Conclusion
We present DPA, an unsupervised domain adaptation method for VLMs that aims at bridging the domain gap between visual and textual representations. DPA introduces a novel idea of dual prototypes, which work as two distinct classifiers, and their outputs are fused via a convex combination. It also ranks pseudo-labels for robust self-training during early training. DPA also enhances the alignment of visual-textual prototypes to improve adaptation performance further. DPA shows improved pseudo-labeling accuracy, which leads to notable performance improvements on various downstream tasks. DPA significantly outperforms zero-shot CLIP and the state-of-the-art methods across 13 downstream vision tasks.